The management of Zigby Manufacturing prepared the following estimated balance sheet for March 2019. ZIGBY MANUFACTURING Estia. Sales for March total 22,800 units. Forecasted sales in units are as follows: April, 22,800; May, 16,000; June, 23,000; anRequired: Prepare the following budgets and other financial information as required. All budgets and other financial informatComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ReqRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Production budRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Raw materialsRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Direct labor bRequired 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Factory overhead budget. (Round perRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Selling expensRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 General and adRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Cash budget. (ZIGBY MANUFACTURING Cash Budget April, May, and June 2019 April May June Beginning cash balance Total cash available Cash payRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Budgeted incomRequired 1 Required 2 Required 3 Required 4 Required 5 Required 6 Required 7 Required 8 Required 9 Required 10 Budgeted balan
Answer #1

Calculation April May June Zigby Manufacturing Sales Budget Req 1 Sales (Units) Selling price per unit Total budgeted sales AMay June Schedule of cash receipts Req 8A Cash sales Credit sales Total cash collections April May 109,440 76,800 437,760 307

1,483,200 1,192,740 290,460 Calculation =61,800*24 =61,800*19.30 Income Statement Req9 Sales Cost of goods sold Gross Profit